DestructoGaming:

Serina wants to solve the following system of equations in the most efficient way. 2 x + 3 y = 18. x + 7 y = 31. She plans to solve for x in the first equation as her first step since both 2 and 3 can be divided into 18. Why is Serina mistaken?

SmokeyBrown:

We can try Serina's strategy for ourselves and see if we run into any issues. So, let's try to solve for x in the first equation [2x + 3y = 18] We can start by getting the 'x' component by itself on one side of the equation; so, let's subtract 3y from both sides to get: [2x = 18 - 3y] Next, if we divide both sides of the equation by 2, we can isolate x: [x = 9 - 3y/2] It looks like we can solve for x, but this did not require us to divide anything by 3, as Serina assumed. Also, we still only have the value of x in terms of y, so there are still steps we would need to take to find the numerical value of x

DestructoGaming:

So.... A?

DestructoGaming:

I've narrowed it down to either A or C.

SmokeyBrown:

@destructogaming wrote:
So.... A?
Oh I didn't even realize it was a multiple choice question, haha. Is it possible you forgot to include that in the original post?

DestructoGaming:

Ok one second.

DestructoGaming:

A. Serina should have solved for x in the second equation because it has a coefficient of 1. B. Serina should have solved for y in the first equation because dividing by 3 instead of by 2 would give a smaller number in the solution. C. Serina should have solved for y in the second equation because it has the largest coefficient. D. Serina should have solved for y in the first equation because the division step will be easier since 18 is divisible by 3.

SmokeyBrown:

It seems like the answer hinges on what it means to solve the equations "in the most efficient way"; I'm not entirely clear on what that means. Out of the options available, I agree that A seems to make the most sense

DestructoGaming:

thanks!

SmokeyBrown:

No prob!